OpenStudy (anonymous):

what is the inverse of f(x)=x^3-1

OpenStudy (anonymous):

Here is a simple thing to remember when doing inverses. Substitute y in for f(x). Then swap your x and y variables. Then solve for y. Example start: f(x) = x³ - 1 y = x³ - 1 x = y³ - 1 Now there are two more steps to solving this, i'll give you a chance to try your hand at finishing :D

OpenStudy (anonymous):

it's cubedrt(x-1), right? if that's the case then how can i make (fof^-1)(x)=x if my function is f(x)=x^3-1 and my inverse is f^-1(x)=cubedrt(x-1)?

OpenStudy (mathstudent55):

You need to take the last equation of @Dragondd320 's excellent explanation and solve for y. How do you solve the following equation for y? \( x = y^3 - 1\) You want to isolate y. What's the first step?